A line through the points (m, -9) and (7, m) has slope m. What is the value of m?
Answer by nerdybill(7384)

A line through the points (m, -9) and (7, m) has slope m. What is the value of m?
.
slope = (y2-y1)/(x2-x1)
m = (m-(-9))/(7-m)
m = (m+9)/(7-m)
m(7-m) = (m+9)
7m-m^2 = m+9
7m = m^2+m+9
0 = m^2-6m+9
0 = (m-3)(m-3)
m = 3